OpenStudy (anonymous):

how can i solve-3r-8=-5r- 12

OpenStudy (anonymous):

add the 5r over 2r-8=-12 add the 8 over 2r=-4 divide by 2 r=-2

jimthompson5910 (jim_thompson5910):

-3r-8=-5r-12 -3r=-5r-12+8 -3r+5r=-12+8 2r=-12+8 2r=-4 r=-4/2 r=-2